# Redundancy-Free View Alignment for Multimodal Human Activity Recognition with Arbitrarily Missing Views

### v2

@@ Abstract
Multimodal multiview learning seeks to integrate information from diverse sources to enhance task performance. Existing approaches often struggle with flexible view configurations, including arbitrary view combinations, numbers of views, and heterogeneous modalities. Focusing on the context of human activity recognition, we propose RALIS, a model that combines multiview contrastive learning with a mixture-of-experts module to support arbitrary view availability during both training and inference. Instead of trying to reconstruct missing views, an adjusted center contrastive loss is used for self-supervised representation learning and view alignment, mitigating the impact of missing views on multiview fusion. This loss formulation allows for the integration of view weights to account for view quality. Additionally, it reduces computational complexity from to , where is the number of views. To address residual discrepancies not captured by contrastive learning, we employ a mixture-of-experts module with a specialized load balancing strategy, tasked with adapting to arbitrary view combinations. We highlight the geometric relationship among components in our model and how they combine well in the latent space. RALIS is validated on four datasets encompassing inertial and human pose modalities, with the number of views ranging from three to nine, demonstrating its performance and flexibility.
@@
multimodal, multiview, missing modality, human activity recognition, contrastive, mixture of experts, redundancy
@@ 1 Introduction
To handle missing views, existing methods often rely on missing-view indicators [12, 45] or reconstruct the missing views using the available ones [37, 35, 18]. However, these approaches supply the model with redundant information because such reconstruction inherently derives from the mutual information among views. In other words, they fill the missing views with information present in the observed views without recovering the underlying information. On the other hand, studies suggest that properly aligning modalities and ensuring coherence among their information can lead to more comprehensive and robust fusion [16]. Multiview contrastive learning pulls different views of the same data sample closer together in the latent space [32]. It is also shown that contrastive learning has a distributional alignment effect [8]. Recent work has integrated contrastive learning into multimodal multiview HAR [7]. Since all sensors observe the same event, they share common information, making this scenario well-suited for contrastive learning, which leverages mutual information to extract robust features across views [3].
@@
To address the above problems, we propose RALIS (Redundancy-free View Alignment with Contrastive Learning and Mixture of Experts), which supports arbitrary view combinations during both training and inference. Each view is first processed by a feature extractor, which can be dedicated to that view or shared among homogeneous views. The resulting features are then combined using an attention-based weighting mechanism. A contrastive loss is used for self-supervised representation learning and to align views, thereby mitigating the impact of missing views on the fusion. This contrastive loss function is trained using the weighted fusion, taking view quality into account. The model head responsible for the main task employs a sparse mixture-of-experts [31] architecture to address the remaining discrepancies among different view combinations. The contributions of this paper are summarized as follows:
We propose RALIS, a multimodal multiview model capable of handling missing views during both training and inference. It shows that view alignment, without missing view reconstruction or filling, can robustly tackle the view missing problem.
Our adjusted center contrastive loss mitigates the impact of view missing on the fusion by aligning views in the hyperspherical latent space. It takes view quality into account and reduces time complexity while still being effective.
@@ 2.3 Multimodal Multiview Learning
Missing modality is a common challenge in multimodal multiview systems. In some cases, views may be intentionally omitted during inference to reduce costs. Several studies have addressed this by leveraging multiple views during training and employing a fixed subset of views for inference, using approaches such as co-training with missing modalities [29] and contrastive learning [14, 27]. To offer greater generality and reduce assumptions about available views, many works address arbitrary missing modalities. For example, the model can be trained to reconstruct missing modalities from the available ones [22, 40, 11]. Some studies fill missing data with trainable embeddings, acting as missing data indicators [45, 12]. Another line of work aligns latent distributions among modalities, then imputes missing modalities with the average of the available ones [35]. However, many methods lack full flexibility: some are limited to homogeneous views, others require particular view combinations in the training set.
3 Avoiding Redundant Information Reconstruction
As mentioned in Sections 1 and 2, existing methods often handle missing views by using missing data indicators, which can be learnable embeddings [12, 45] or zero/random values [38]. This approach informs the model which data samples have missing views by filling those missing entries with indicators, but it does not provide any additional information beyond indicating absence. Many other methods reconstruct the missing views using the available ones [22, 40, 11]. Although this approach seems to recover the missing data, the reconstruction fundamentally depends on the information already present in the observed views. In other words, the reconstructed data inherently derives from the mutual information among views.
where
is the intrinsic entropy of determined by its marginal distribution , and is not changed by the reconstruction model. Therefore, the model is essentially learning to maximize the mutual information . If we directly train the model to maximize this mutual information, it can achieve the same effect without constructing redundant data, thereby reducing workload. Prior studies have shown that multiview contrastive learning can maximize mutual information among views [3, 32], making it suitable for this task.
@@ 4 Proposed Method
Figure 1 illustrates the RALIS model. The input data from each view is passed through its corresponding encoder to extract feature vectors, which are then normalized to the same magnitude. An attention module assigns a weight to each view within a data sample based on task relevance, enabling view fusion via a weighted sum. The fused vector also goes through magnitude normalization, putting it onto the same hyperspherical feature space as the individual views. During training, both individual and fused representations are passed into the MoE head for classification and are also used for contrastive learning. This ensures the model learns both useful view-specific information and mutual information among views.
@@ 4.1 Adjusted Center Contrastive Loss
Besides the discussion in Section 3, we use contrastive loss in RALIS for two reasons. First, contrastive loss pulls similar data points (positives) together and pushes dissimilar ones (negatives) apart [36]. In this multiview setting, we consider views of the same data sample as positives and views from different samples in a training batch as negatives. Since we fuse views by finding their center in the feature space, the number and distribution of views influence the fusion. By pulling views closer together, contrastive loss improves the fusion’s robustness to view missing (Figure 2).
Second, contrastive learning is widely used for self-supervised representation learning [20, 7]. Even without fusion, studies have shown that multiview contrastive learning can improve uniview models [14, 27].
@@ 4.1.1 Loss Function Definition
To compute multiview contrastive loss, the full graph approach contrasts all pairs, pulling all views together (i.e., indirectly pulling all views toward a center). The core view approach contrasts each view with a designated core view, aligning all views toward this core [32]. Full graph often yields better performance than core view, as it does not rely on a single view [17]. However, the core view approach is more efficient, requiring only pairs compared to pairs for the full graph. Our method instead aligns all views directly to the center by contrasting each view with the center of the other views (Figure 3). This reduces time complexity while preserving the benefits of the full graph approach. Additionally, view weights can be integrated straightforwardly to adjust the center, enabling control over the influence of individual views with varying quality. Conversely, a full graph approach would require assigning a pairwise weight to every view pair.
@@
The weight term puts the center closer to higher-quality views, causing all views to be pulled more toward those of higher quality. The term acts as a loss weight, assigning a lower weight when the loss function tries to pull a high-quality view (i.e., when is large) toward others. This weight will be discussed in Section 4.2. The loss term is divided by instead of to compensate for the scale decrease caused by the weight . Specifically, the total scaling factor is:
@@ 4.3.1 MoE Head
For each input of an individual view or the fusion, the gate outputs a set of weights across the experts. Only the top experts are kept and their weights are passed into a softmax function, while the rest are assigned a weight of 0.
@@ 4.3.2 Load Balancing
4.4 RALIS in A Hyperspherical Feature Space
In this section, we look at the geometric properties of RALIS’s components and their compatibility in a hyperspherical latent space.
@@ 5.1.2 Baselines
The following methods are included in the comparison: CMC [32], COCOA [9], Flex-MoE [45], FuseMoE [12], and ShaSpec [35]. CMC and COCOA are not explicitly designed for missing views, but since they use contrastive learning and process views independently, the same missing view handling strategy as in RALIS can be applied (Figure 2); these methods are marked with an asterisk in the result tables. Flex-MoE and FuseMoE both use learnable embeddings as missing view indicators and employ MoE to accommodate arbitrary view combinations. ShaSpec learns separate representations for view-specific and shared information; any missing view is reconstructed from the shared features of the available views.
@@ 5.2.2 Missing Training Data
Table 4 shows that RALIS achieves the highest scores overall. The performance gap between RALIS and other methods becomes more pronounced, particularly on the DailySport and RealDisp datasets. This demonstrates RALIS’s robustness to missing views compared to the baselines.
@@ 5.4 More Results with Random View Missing Rate
Table 6 shows the results for the baseline models, ablation study, and our proposed model. Overall, RALIS achieves the best scores. With a different view missing pattern, these results still remain consistent with the previous experiments.
@@ 7 Conclusion
This paper introduces RALIS for flexible multimodal multiview HAR. RALIS can handle arbitrarily missing views during both training and inference, while avoiding redundant missing view reconstruction.

### v3

# Align and Adapt: Multimodal Multiview Human Activity Recognition under Arbitrary View Combinations

@@ Abstract
Multimodal multiview learning seeks to integrate information from diverse sources to enhance task performance. Existing approaches often struggle with flexible view configurations, including arbitrary view combinations, numbers of views, and heterogeneous modalities. Focusing on the context of human activity recognition, we propose AliAd, a model that combines multiview contrastive learning with a mixture-of-experts module to support arbitrary view availability during both training and inference. Instead of trying to reconstruct missing views, an adjusted center contrastive loss is used for self-supervised representation learning and view alignment, mitigating the impact of missing views on multiview fusion. This loss formulation allows for the integration of view weights to account for view quality. Additionally, it reduces computational complexity from to , where is the number of views. To address residual discrepancies not captured by contrastive learning, we employ a mixture-of-experts module with a specialized load balancing strategy, tasked with adapting to arbitrary view combinations. We highlight the geometric relationship among components in our model and how they combine well in the latent space. AliAd is validated on four datasets encompassing inertial and human pose modalities, with the number of views ranging from three to nine, demonstrating its performance and flexibility.
@@
contrastive, human activity recognition, missing modality, mixture of experts, multimodal, multiview
@@ 1 Introduction
To handle missing views, existing methods often rely on missing-view indicators [12, 46] or reconstruct the missing views using the available ones [38, 36, 18]. However, such reconstruction inherently derives from the mutual information among views. In other words, they fill the missing views with information present in the observed views without recovering the actual missing information. Furthermore, training a separate reconstruction model for each view becomes impractical as the number of views increases. On the other hand, studies suggest that properly aligning modalities and ensuring coherence among their information can lead to more comprehensive and robust fusion [16]. Multiview contrastive learning pulls different views of the same data sample closer together in the latent space [33]. It is also shown that contrastive learning has a distributional alignment effect [8]. Recent work has integrated contrastive learning into multimodal multiview HAR [7]. Since all sensors observe the same event, they share common information, making this scenario well-suited for contrastive learning, which leverages mutual information to extract robust features across views [3].
@@
To address the above problems, we propose AliAd (Align and Adapt), which supports arbitrary view combinations during both training and inference. Each view is first processed by a feature extractor, which can be dedicated to that view or shared among homogeneous views. The resulting features are then combined using an attention-based weighting mechanism. A contrastive loss is used for self-supervised representation learning and to align views, thereby mitigating the impact of missing views on the fusion. This contrastive loss function also takes view quality into account. The model head responsible for the main task employs a sparse mixture-of-experts [32] architecture to address the remaining discrepancies among different view combinations. The contributions of this paper are summarized as follows:
We propose AliAd, a multimodal multiview model capable of handling missing views during both training and inference. It shows that view alignment, without missing view reconstruction or filling, can robustly tackle the view missing problem.
Our adjusted center contrastive loss mitigates the impact of view missing on the fusion by aligning views in the hyperspherical latent space. It takes view quality into account and reduces time complexity.
@@ 2.3 Multimodal Multiview Learning
Missing modality is a common challenge in multimodal multiview systems. In some cases, views may be intentionally omitted during inference to reduce costs. Several studies have addressed this by leveraging multiple views during training and employing a fixed subset of views for inference, using approaches such as co-training with missing modalities [30] and contrastive learning [14, 27]. To offer greater generality and reduce assumptions about available views, many works address arbitrary missing modalities. For example, the model can be trained to reconstruct missing modalities from the available ones [22, 11]. Some studies employ modality alignment techniques; however, these methods still require missing modality reconstruction afterwards [29, 41]. Some studies fill missing data with trainable embeddings, acting as missing data indicators [46, 12]. Another line of work aligns latent distributions among modalities, then imputes missing modalities with the average of the available ones [36]. However, many methods lack full flexibility: some are limited to homogeneous views, others require particular view combinations in the training set.
@@ 3 Proposed Method
Figure 1 illustrates the AliAd model. The input data from each view is passed through its corresponding encoder to extract feature vectors, which are then normalized to the same magnitude. An attention module assigns a weight to each view within a data sample based on task relevance, enabling view fusion via a weighted sum. The fused vector also goes through magnitude normalization, putting it onto the same hypersphere as the individual views. During training, both individual and fused representations are passed into the MoE head for classification and are also used for contrastive learning. This ensures the model learns both useful view-specific information and mutual information among views.
@@ 3.1 Adjusted Center Contrastive Loss
3.1.1 Why Contrastive Loss?
Existing methods often handle missing views by using missing data indicators, which can be learnable embeddings [12, 46] or zero/random values [39]. This approach informs the model which data samples have missing views by filling those missing entries with indicators, but it does not provide any additional information beyond indicating absence. Many other methods reconstruct the missing views using the available ones [22, 29, 41, 11]. Although this approach seems to recover the missing data, the reconstruction fundamentally depends on the information already present in the observed views. In other words, the reconstructed data inherently derives from the mutual information among views. Furthermore, training a reconstruction model for each view is less practical when the number of views gets large.
where is the intrinsic entropy of determined by its marginal distribution , and is not changed by the reconstruction model. Therefore, the model is essentially learning to maximize the mutual information . If we directly train the model to maximize this mutual information, it can achieve the same effect without constructing redundant data, thereby reducing workload. Prior studies have shown that multiview contrastive learning can maximize mutual information among views [3, 33], making it suitable for this task.
We also use contrastive loss for two other reasons. First, contrastive loss pulls similar data points (positives) together and pushes dissimilar ones (negatives) apart [37]. In this multiview setting, we consider views of the same data sample as positives and views from different samples in a training batch as negatives. Since we fuse views by finding their center in the feature space, the number and distribution of views influence the fusion. By pulling views closer together, contrastive loss improves the fusion’s robustness to view missing (Figure 2).
@@ 3.1.2 Loss Function Definition
To compute multiview contrastive loss, the full graph approach contrasts all pairs, indirectly pulling all views together. The core view approach contrasts each view with a designated core view, aligning all views toward this core [33]. Full graph often yields better results than core view, as it does not rely on a single view [17]. However, core view is more efficient, requiring only pairs compared to pairs in a full graph. Our method instead aligns all views directly to the center by contrasting each view with the center of the other views (Figure 3). This reduces time complexity while preserving the benefits of the full graph approach. Also, view weights can be integrated straightforwardly to adjust the center, enabling control over the influence of individual views with varying quality. Conversely, a full graph approach would require assigning a pairwise weight to every view pair.
@@
The weight term puts the center closer to higher-quality views, causing all views to be pulled more toward those of higher quality. The term acts as a loss weight, assigning a lower weight when the loss function tries to pull a high-quality view toward others. This weight will be discussed in Section 3.2. The loss term is divided by instead of to compensate for the scale decrease caused by the weight . Specifically, the total scaling factor is:
@@ 3.3.1 MoE Head
For each input of an individual view or the fusion, the gate outputs a set of weights across the experts. Only the top experts are kept, and their weights are passed into a softmax function, while the rest are assigned a weight of 0.
@@ 3.3.2 Load Balancing
3.4 AliAd in A Hyperspherical Feature Space
In this section, we look at the geometric properties of AliAd’s components and their compatibility in a hyperspherical latent space.
@@ 4.1.2 Baselines
The following methods are included in the comparison: CMC [33], COCOA [9], Flex-MoE [46], FuseMoE [12], and ShaSpec [36]. CMC and COCOA are not explicitly designed for missing views, but since they use contrastive learning and process views independently, missing views are naturally handled the same way as in our method (Figure 2). Flex-MoE and FuseMoE both use learnable embeddings as missing view indicators and employ MoE to accommodate arbitrary view combinations. They also use the transformer architecture to model cross-view interactions. ShaSpec learns separate representations for view-specific and shared information; any missing view is reconstructed from the shared features of the available views.
@@ 4.2.2 Missing Training Data
Table 4 shows that AliAd achieves the highest scores overall. The performance gap between AliAd and other methods becomes more pronounced, particularly on the DailySport and RealDisp datasets. This demonstrates AliAd’s robustness to missing views compared to the baselines.
@@ 4.4 More Results with Random View Missing Rate
Table 6 shows the results for the baseline models, ablation study, and our proposed model. Overall, AliAd achieves the best scores. With a different view missing pattern, these results still remain consistent with the previous experiments.
@@ 6 Conclusion
This paper introduces AliAd for flexible multimodal multiview HAR. AliAd can handle arbitrarily missing views during both training and inference, while avoiding unnecessary data reconstruction.